Ikerd Company applies manufacturing overhead to jobs on the basis of machine hours used. Overhead costs are estimated to total $317,250 for the year, and machine usage is estimated at 126,900 hours. For the year, $334,900 of overhead costs are incurred and 130,000 hours are used. Incorrect answer. Your answer is incorrect. Try again. Compute the manufacturing overhead rate for the year. (Round answer to 2 decimal places, e.g. 1.25.) Manufacturing overhead rate $Entry field with incorrect answer per machine hour SHOW LIST OF ACCOUNTS LINK TO TEXT LINK TO TEXT Incorrect answer. Your answer is incorrect. Try again. What is the amount of under- or overapplied overhead at December 31? Manufacturing Overhead $Entry field with incorrect answer Entry field with incorrect answer now contains modified data SHOW LIST OF ACCOUNTS LINK TO TEXT LINK TO TEXT Partially correct answer. Your answer is partially correct. Try again. Prepare the adjusting entry to assign the under- or overapplied overhead for the year to cost of goods sold. (Credit account titles are automatically indented when amount is entered. Do not indent manually.) Account Titles and Explanation Debit Credit Entry field with incorrect answer Entry field with incorrect answer now contains modified data Entry field with correct answer Entry field with incorrect answer Entry field with incorrect answer now contains modified data Entry field with incorrect answer now contains modified data Click if you would like to Show Work for this question: Open Show Work
Answer #1
1
Manufacturing overhead rate 2.50 per machine hour =317250/126900
2
Manufacturing overhead applied 325000 =130000*2.50
Manufacturing overhead 9900 Underapplied =334900-325000
3
Debit Credit
Cost of goods sold 9900
      Manufacturing overhead 9900